A manufacturer of graphing calculators has determined that 9,000 calculators per week will be sold at a price of $85 per calculator. At a price of $80, it is estimated that 11,000 calculators will be sold.

1. Determine a linear function that predicts the number of calculators that will be sold per week at a price of x dollars.


2. Use the function to find the cost of using the cellular phone for 13 min in 1 month.
1
Expert's answer
2020-09-30T19:35:11-0400

Solution.

A linear function has the following form:

"y=a+b\cdot x" , where

y - number of calculators that will be sold per week (pieces/week);

x - calculator price ($);

a - constant (pieces/week);

b - constant (pieces/(week⋅$)).

Find constant b by subtraction of formulas:

"9000=a+b\cdot 85" ;

"11000=a+b\cdot 80" , where

"9000-11000=a+b\cdot 85-(a+b\cdot 80)" ;

"-2000=b\cdot 5" ;

"b=-400" (pieces/(week⋅$)).

Find constant a:

"9000=a-400\cdot 85" ;

"a=9000+400\cdot 85=43000" (pieces/week).

Answer 1: "y=43000-400\cdot x".

Answer 2: We can't use upper function, because constants must have other dimensions (a must have (min/month) dimension; b must have (min/(month⋅$)) dimension). Accordingly we can use upper function to find only the number of calculators that will be sold per week at a price of x dollars.